CH/Doc/98/01

Minutes

Charissa Planning Meeting

Meeting held on: 11 September 1998

Meeting held at: Staffordshire

Present: BRF,WNC,RC,SPGC,NMC,RLC,NC,NJD,MF,RAH,RCL,JTM,

WDMR,SMS,RPW,DDW

Distribution: above,char-meeting@nnga.dl.ac.uk by e-mail/file-server

Charissa two-day meeting

B.R. Fulton

Thanks to all those who prepared talks for the first day, which was a very useful update on progress with existing experiments. I have a copy of all the overheads and duplicates will be posted out to everyone next week. A big thanks also to Roger for organising everything, including the excellent evening event for Dick’s retirement. On the second day we discussed the grant submission and agreed that we would go for a June siubmission for 4 years. We also agreed that our first step is to produce drafts on the main physics themes and on the major equipment development ideas. These are to be ready for the end of November. A meeting will be arranged for mid December to consider how these can be merged together and to agree on the next steps in preparing the grant submission. The table summarises agreed actions:

Physics Theme

1. Molecular like states in exotic nuclei:

(a) Radioactive beams SPGC/MF

(b) Stable beams MF/SPGC/RLC

2. Spectroscopy of exotic nuclei with transfer reactions WNC/RC/RCL

3. C+C and similar systems WDMR/NJD

4. Reactions with losely bound nuclei BRF/RPW Equipment development

1. Mass resolution in MEGHA NMC/MF/SPGC

2. New array at GANIL WNC/RC/RL

3. Possible new beamline and chamber at GANIL MF

In addition, on the immediate problem of the PSD detector failures:

1. Beta source tests to see if caused by electrons NMC/BRF

2. In beam tests (at Strasbourg?)

MF/RLC

W.N. Catford, Secretary September 21, 1998